Question #2 Consider the general demand function and the supply function for Dates (Tamr) in (tons) below: la = 60 - 2P + 0.02- Are dates normal good or inferior good? Explain?3- What are the determinants of demand in the general demand function above?4- Are dates substitute with the related product (chocolates)? Explain?5- What is the effect of the increased quantity demanded by 20 tons, on the equilibrium price and quantity demanded and suppl
Answer #1

Question 2 Considering the queral dewal fruction and the Supply for ction for Dabs (Tan ) in (tous) below! Qd = 60-2P toolmtosupply Now, Qs = 600+2P =) Qs - 600 = 2p. . O.S QS – 300 =p a) p= 0.5Q5 -3.00 The rola i the a relationship betwan quantity s4 Now, - Qd = 60-ap+ 0.01m tos po tole demand between Now dates cross-price elasticity of and chocolates will be Ec= sado por